# Quasi-randomization tests for network interference

### v1

@@ 4.3   Test statistic
The choice of the test statistic is not unique, and the validity of the proposed test is upheld for any choice of test statistic. Test statistics capture departures from the null distribution, and better-suited test statistics in responding to deviations from the null will improve the power of the test. Here, we propose a variation of the Has-Treated-Neighbour test statistic considered in Athey et al. (2018). This looks at the difference in average treatment effect amongst control units with at least one treated neighbor and control units with only control units as neighbors. Essentially, we treat focal units as control units and auxiliary units as treated units.
$T_{I_{c}}(G|Z_{obs})$ $=$ (4)
$\frac{\sum_{i=1}^{N}Y_{i}\cdot\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}>0,Z_{i_{obs}}=0)}{\sum_{i=1}^{N}\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}>0,Z_{i_{obs}}=0)}-\frac{\sum_{i=1}^{N}Y_{i}\cdot\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}=0,Z_{i_{obs}}=0)}{\sum_{i=1}^{N}\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}=0,Z_{i_{obs}}=0)},$
Note that the above test statistic considers the spillover on the control units from the treated. Under the null of no spillover effect for all the units regardless of direct treatment status, one can similarly define the above test statistic for the treated units.
$T_{I_{t}}(G|Z_{obs})$ $=$ (5)
$\frac{\sum_{i=1}^{N}Y_{i}\cdot\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}>0,Z_{i_{obs}}=1)}{\sum_{i=1}^{N}\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}>0,Z_{i_{obs}}=1)}-\frac{\sum_{i=1}^{N}Y_{i}\cdot\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}=0,Z_{i_{obs}}=1)}{\sum_{i=1}^{N}\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}=0,Z_{i_{obs}}=1)},$
@@
$T_{I}(G|Z_{obs})=\frac{N_{c}}{N}\cdot T_{I_{c}}+\frac{N_{t}}{N}\cdot T_{I_{t}}$ (6)
It should be noted the above test statistics can only be defined if the sub-populations in consideration are non-empty. For example, if there exists no control unit with all neighbors in control or the number of such units is few, we deal with either a definition or power issue. Thus, we propose a generalization of the above-stated test statistic for robustness. We look at the difference in average treatment effect among control units with greater than 0.75 quantile of the proportion of treated units and control units with less than 0.25 quantile of the proportion of treated units.
$T_{quant_{c}}(G|Z_{obs})$ $=$ (7)
$\frac{\sum_{i=1}^{N}Y_{i}\cdot\mathcal{I}(Z_{i_{obs}}=0)\cdot\mathcal{I}(\frac{\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}}{\sum_{j=1}^{N}G_{ij}}\geq p_{0.75})}{|P_{0.75}|}-\frac{\sum_{i=1}^{N}Y_{i}\cdot\mathcal{I}(Z_{i_{obs}}=0)\cdot\mathcal{I}(\frac{\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}}{\sum_{j=1}^{N}G_{ij}}\leq p_{0.25})}{|P_{0.25}|},$
@@
$P_{0.25}$ and $P_{0.75}$ represent the set of elements in $P$ in the first and third quartile, respectively. In a similar fashion to the test statistic $T_{I}(G|Z_{obs})$, we define the above for treated units and take a weighted average for the two test statistic to obtain $T_{quant}(G|Z_{obs})$. We remark that one can improve the test statistic with the knowledge of appropriate covariates or nodal attributes. This can be incorporated by considering regression-adjusted test statistics taking residuals for the potential outcome (Rosenbaum (2002)). One can also build test statistics based on models for interference structures. For example, parametric test statistics can be constructed of the linear-in-means model of interference proposed by Manski (1993). See Athey et al. (2018) for a more elaborate discussion on this.
@@ 5   Simulation study
We present Monte Carlo simulations to validate our proposed testing procedure. We consider the test statistic based on the quantile of the proportion of treated neighbors in the simulations, as discussed in Section 4.3. We will describe the data-generating process and the general setup in the next section. We conduct the simulation study for two experimental designs, namely, completely randomized experiments and cluster randomized experiments. As a robustness check, we repeat the study for two network-generating processes.
Table 1: Rejection rates for the null hypothesis of no spillover effects for completely randomized experiment
Network Distribution Direct Effect Spillover Effect Degree Permutation Test Conditional Focal Test
$T_{quant}$ $T_{elc}$
Small world network 0 0 0.049 0.050
4 0 0.046 0.046
0 0.4 0.106 0.112222We reported the number presented in Athey et al. (2018). We obtained a rejection rate of 0.073 in our replication.
4 0.4 0.101 0.060
Stochastic block model 0 0 0.062 0.054
4 0 0.055 0.052
0 0.4 0.130 0.065
4 0.4 0.140 0.048
@@ 5.1   Data generating mechanism
We note that $\epsilon_{i}$ can be viewed as the baseline level of outcomes for all the units if they were untreated. Without loss of generality, we take the mean of the normal distribution to be zero, but it can be arbitrarily chosen. We first present the setup for a completely randomized experiment. The treatment assignment vector is chosen by randomly sampling $N_{t}$ units out of $N$ units and assigning them treatment (that is, $Z_{i}=1$ for all $i$ corresponding to units in the random sample). We state that $N_{t}$ is fixed and pre-determined by the experimenter. In the performed simulations, we have $599$ units and $300$ treated units in the model in one case. We study the setup for two generating processes, namely, small-world networks and the stochastic block model. Both models exhibit a high degree of clustering, which is characterized by the clustering coefficient. The clustering coefficient is the number of connections in a randomly chosen node and its neighborhood relative to the total number of connections possible in the neighborhood. Consider a regular graph, which is a graph where all the nodes have the same number of connections. Let us denote the number of connections a unit has in a regular graph by $K$. Such graphs have a high clustering coefficient but also distance between two nodes in such graphs is also high. It is observed in real-life networks that the average distance between two nodes is low (Amaral et al. (2000)). To obtain a graph with a high degree of clustering but a low average path length, one can rewire the edges of a $K$-regular graph with some probability, say $p$. This parametrization characterizes a small world network. The higher the probability of rewiring, the lower the average path length gets. Another class of random graph models that achieves the same property of a high clustering coefficient but a low average path length is that of the stochastic block model. Consider a graph of size $N$ with a partition of nodes. The size of the parts are $\{n_{1},n_{2},...,n_{K}\}$. These parts represent clusters in the graph, and we can model a random graph using it by defining the probability of an edge between nodes within a part and between two parts. Hence, we obtain a probability (symmetric) matrix of size $K\times K$. Note that because we want a high degree of clustering in the graph, the diagonal elements of the probability matrix are typically higher than the off-diagonal elements. The probability matrix is also referred to as the preference matrix in the literature. We now define the setup for the simulation study:
@@
    Direct treatment effect: We take $\tau_{direct}$ equal to 4 and 0 in our simulations.
@@
    Test statistic: We consider the $T_{quant}$ test statistic described in Section 4.3. It takes the difference in the treatment effect of control units in the first quartile of the proportion of treated neighbors to control units in the last quartile of the proportion of treated neighbors. We define the same for treated units and take the weighted average of both.
The null randomization distribution was generated by sampling graphs with the same degree sequence as the observed graph. The null distribution was approximated by generating 1000 graphs with the same degree sequence as the observed graph. The p-value is calculated as the sample proportion of times the imputed test statistic for different graph realizations, with the same degree sequence as the original graph, is more than the observed test statistic. If the p-value obtained is less than the chosen significance level of 0.05, we reject the null hypothesis. We replicated the above-described setup 4000 times and calculated the proportion of times we rejected the null. We present this result, namely rejection rate, in the Table 1. For robustness, we perform the simulations for two network-generating processes. We obtain the rejection rates for the same setup for the procedure proposed by Athey et al. (2018). We implement random focal vertex selection and show our results for the Edge-Level-Contrast test statistic (Athey et al. (2018)).
We then perform the simulations for cluster randomized experimental setup. We present the results in Table 2. The total number of units in the simulation is the same as before, that is, 599. We obtain the clusters in the graph by implementing the epsilon-net clustering algorithm, described in Algorithm 1 (Appendix A). We take epsilon to be equal to 3. This implies all the clusters contain units that are less than 3 hops away from each other, whereas between cluster distance is at least 3 hops for any two units. We take the probability of treating a cluster to be 0.5. We implemented a Bernoulli trial at the cluster level. The null distribution is again approximated by taking 1000 samples of graphs with the same degree sequence as the original graph. We replicated this setup 2000 times and calculated the proportion of times the proposed test rejected the null. We reduce the replication size to 2000 for cluster randomized setup observing appropriate convergence of the p-value estimate. This was performed for two network-generating processes as before.
@@
Network Distribution Direct Effect Spillover Effect Degree Permutation Test
$T_{quant}$
Small world network 0 0 0.055
4 0 0.055
0 0.4 0.213
4 0.4 0.217
Stochastic block model 0 0 0.053
4 0 0.055
0 0.4 0.168
4 0.4 0.164
@@ 5.2   Comparison with related work
We see that our proposed method has type 1 error controlled appropriately. This corresponds to rows with $\tau_{spill}$ equal to zero in Tables 1 and 2. When $\tau_{spill}$ is equal to 0.4, we see that our method performs with competitive power compared to existing methods in the literature (Athey et al. (2018)). We highlight the case of $\tau_{direct}$ equal to 4 and $\tau_{spill}$ equal to 0.4, where our method displays substantial improvement. The significance of the setup lies in its resemblance to many naturally occurring settings, where the direct effect is meaningfully more than the spillover effect (for example, see Cai et al. (2015)). We see that when the clustering of the network-generating process increases, as is the case with the stochastic block model, our proposed method displays a boost in power. This is in contrast with the method proposed by Athey et al. (2018), where the method does not retain power. An interesting result is in the case of cluster-randomized experiments where a higher degree of clustering shows a drop in power. This can be observed when comparing the power values of the small-world network to the stochastic block model. This can be attributed to a reduction of sample space of imputed values of the test statistic. Hence, a higher degree of clustering increases the cluster sizes, which restricts the direct treatment assignment. The observed graph in the case of the small-world network had 16 clusters, as compared to 11 clusters in the case of the stochastic block model. We also perform the simulation study for cluster randomized experiments for the methods proposed in Athey et al. (2018) and Puelz et al. (2022). As discussed in Section 3, we encounter the degenerate null distribution and hence, do not present the corresponding results. This is more prominent in the method proposed by Athey et al. (2018), in comparison with Puelz et al. (2022). Since Puelz et al. (2022) takes into account the design of the experiment when selecting focal units, we see more robustness of the biclique test (Puelz et al. (2022)) where degenerate null distributions are reduced in frequency. Overall, our method stands more robust to degeneracy and performs well in terms of power in our simulations.
6   Application: Weather insurance adoption in rural China
@@
We re-analyze a field experiment conducted in rural China on rice crop farmers studying neighborhood interference effects (Cai et al. (2015)). The research motive behind the experiment was to see if social networks among the farmers have any influence on weather insurance adoption. The experiment involved conducting intensive training sessions to promote crop insurance products to farmers against extreme weather. A primary goal of the experiment was to identify if there is information dissemination in the farmers’ community and, if so, what mechanisms drive these information exchanges. These are important questions to be answered in this setting as they are consequential to government policy action. A deeper understanding of insurance product adoption via social network mechanisms helps devise a more cost-effective government policy campaign to enhance product adoption among the target community. Once interaction mechanisms can be established, better “seeding” or targeting strategies can be rolled out for future campaigns (Ok et al. (2014)). Such strategies identify influential figures in the community to maximize the interference effect and, hence, the program’s overall impact. Another important reason to study interference in such a setup is to avoid biased estimates and accurately evaluate policy performance measures.
@@
We discuss the experimental design setup by the researchers. The hypothesis of interest is the underlying assumption that access to better information about weather insurance products would increase the adoption of the product in the rice farmers’ community. To capture access to information differentially as treatment and control, two types of information sessions were created: simple sessions of 20 minutes, which provided an introduction to the insurance contract, and intensive sessions of 45 minutes that covered the insurance contract in much more detail, including explaining to them what the benefits are of taking up insurance. To capture information dissemination across the social network, these sessions were conducted in two rounds with a gap of 3 days. The idea behind this setup is to allow the farmers some time to discuss the insurance product among themselves. The interference effect is identified by examining whether households in round 2 with more friends who received intensive sessions in round 1 displayed an enhanced insurance take-up rate. Second-round participants were further randomized into three groups. After the round 2 sessions, the first group was given no additional information, the second group was told about the overall take-up rate in the previous round, and the third group was told in detail about who purchased the product and who did not. This additional information randomization of the second group was done to derive the main mechanism via which information diffusion took place. For illustration purposes, we do not analyze this part of the experiment.
@@
There were a total of 4902 households across 47 different villages in the experiment. The social network among a village is known and is assumed to be independent between two distinct villages. Each household is randomized among the 4 different groups: (i) Receiving simple session in Round 1,(ii) Receiving intensive session in Round 1, (iii) Receiving simple session in Round 2, (iv) Receiving intensive session in Round 2. After the session ends, the households make the decision to purchase insurance. This is captured as a binary outcome variable of the model, with 1 indicating a purchase decision made and 0 when no purchase is made. Under the assumption that no information diffusion can take place between participant households of Round 1 because of the immediate purchase decision the household had to make, we can assume there are no edges among the household participants in Round 1.
@@
In our analysis, we take all the households receiving simple sessions under the control group and intensive sessions under the treated group. Since spillover can only occur from the treated group of Round 1 to the control group of Round 2, we only consider the network structure amongst these two groups in our study. We retain the network structure among the households between the two rounds and nullify the remaining network structure. This includes network structure among households in Round 1 only and network structure among households in Round 2 only. Our analysis was run on 433 households of randomly chosen villages in the experiment, of which 222 were in the treated group. We ran our proposed Procedure 2 and obtained a p-value of 0.0002, which indicates the presence of interference.

### v2

@@ 4.3 Test statistic
The choice of the test statistic is not unique, and the validity of the proposed test is upheld for any choice of test statistic. Test statistics capture departures from the null distribution, and better-suited test statistics in responding to deviations from the null will improve the power of the test. Here, we propose a variation of the Has-Treated-Neighbour test statistic considered in Athey et al. (2018). This looks at the difference in average treatment effect amongst control units with at least one treated neighbor and control units with only control units as neighbors. Let $T_{I_{c}}(G|Z_{obs})=$
$\frac{\sum_{i=1}^{N}Y_{i}\cdot\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}>0,Z_{i_{obs}}=0)}{\sum_{i=1}^{N}\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}>0,Z_{i_{obs}}=0)}-\frac{\sum_{i=1}^{N}Y_{i}\cdot\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}=0,Z_{i_{obs}}=0)}{\sum_{i=1}^{N}\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}=0,Z_{i_{obs}}=0)}.$ (3)
The above test statistic considers the spillover on the control units from the treated. Under the null of no spillover effect for all the units regardless of direct treatment status, one can also define the above test statistic for the treated units. Similarly, let $T_{I_{t}}(G|Z_{obs})$ equals
$\frac{\sum_{i=1}^{N}Y_{i}\cdot\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}>0,Z_{i_{obs}}=1)}{\sum_{i=1}^{N}\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}>0,Z_{i_{obs}}=1)}-\frac{\sum_{i=1}^{N}Y_{i}\cdot\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}=0,Z_{i_{obs}}=1)}{\sum_{i=1}^{N}\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}=0,Z_{i_{obs}}=1)}.$ (4)
@@
$T_{I}(G|Z_{obs})=\frac{N_{c}}{N}\cdot T_{I_{c}}+\frac{N_{t}}{N}\cdot T_{I_{t}}.$ (5)
It should be noted the above test statistics can only be defined if the sub-populations in consideration are non-empty. For example, if no control unit exists with all neighbors in control or the number of such units is few, we deal with either a definition or power issue. Thus, we propose a generalization of the above-stated test statistic for robustness. We look at the difference in average treatment effect among control units with greater than 0.75 quantile of the proportion of treated units and control units with less than 0.25 quantile of the proportion of treated units. Let $T_{quant_{c}}(G|Z_{obs})=$
$\frac{\sum_{i=1}^{N}Y_{i}\cdot\mathcal{I}(Z_{i_{obs}}=0)\cdot\mathcal{I}(\frac{\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}}{\sum_{j=1}^{N}G_{ij}}\geq p_{0.75})}{|P_{0.75}|}$ (6)
$-\frac{\sum_{i=1}^{N}Y_{i}\cdot\mathcal{I}(Z_{i_{obs}}=0)\cdot\mathcal{I}(\frac{\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}}{\sum_{j=1}^{N}G_{ij}}\leq p_{0.25})}{|P_{0.25}|}.$
@@
$P_{0.25}$ and $P_{0.75}$ represent the set of elements in $P$ in the first and third quartile, respectively. In a similar fashion to the test statistic $T_{I}(G|Z_{obs})$, we define the above for treated units and take a weighted average for the two test statistics to obtain $T_{quant}(G|Z_{obs})$.
We remark that one can improve the test statistic with the knowledge of appropriate covariates or nodal attributes. This can be incorporated by considering regression-adjusted test statistics taking residuals for the potential outcome (Rosenbaum (2002)). One can also build test statistics based on models for interference structures. For example, parametric test statistics can be constructed of the linear-in-means model of interference proposed by Manski (1993). We also consider more test statistics, given that the choice is not unique and different test statistics would differ in their power. We take the test statistic proposed in Bond et al. (2012), which takes the difference between the average over all the edges where the neighbor is a treated or control unit.
@@ 5 Simulation study
We present Monte Carlo simulations to validate our proposed testing procedure. We consider two test statistics, one that is based on the quantile of the proportion of treated neighbors in the simulations and another that is weighted by treated neighbors (first proposed in Bond et al. (2012)), as discussed in Section 4.3. The following subsection will describe the data-generating process and the general setup. We conduct the simulation study for two experimental designs, namely, completely randomized experiments and cluster randomized experiments. As a robustness check, we repeat the study for two network-generating processes.
@@ 5.1 Data generating mechanism
We note that $\epsilon_{i}$ can be viewed as the baseline level of outcomes for all the units if they were untreated. Without loss of generality, we take the mean of the normal distribution to be zero, but it can be arbitrarily chosen. We first present the setup for a completely randomized experiment. The treatment assignment vector is chosen by randomly sampling $N_{t}$ units out of $N$ units and assigning them treatment (that is, $Z_{i}=1$ for all $i$ corresponding to units in the random sample). We state that $N_{t}$ is fixed and pre-determined by the experimenter. In the performed simulations, we have $599$ total units and $300$ treated units in the model in one case. We study the setup for two generating processes: small-world networks and the stochastic block model. Both models exhibit a high degree of clustering, characterized by the clustering coefficient. The clustering coefficient is the number of connections in a randomly chosen node and its neighborhood relative to the total number of connections possible. Consider a regular graph where all the nodes have the same cardinality of connections. Let us denote the number of connections a unit has in a regular graph by $K$. Such graphs have a high clustering coefficient, but the distance between two nodes in such graphs is also high.
Table 1: Rejection rates for the null hypothesis of no spillover effects for completely randomized experiment.
Network Distribution Direct Effect Spillover Effect Degree Permutation Test Isomorphism Permutation Test Conditional Focal Test
$T_{bond}$ $T_{quant}$ $T_{bond}$ $T_{quant}$ $T_{elc}$
Small world network 0 0 0.045 0.049 0.049 0.047 0.050
4 0 0.054 0.046 0.056 0.045 0.046
0 0.4 0.187 0.148 0.191 0.162 0.112111We reported the number presented in Athey et al. (2018). We obtained a rejection rate of 0.073 in our replication.
4 0.4 0.074 0.151 0.064 0.147 0.060222We obtain the same rejection rate in our replication as reported in Athey et al. (2015).
Stochastic block model 0 0 0.045 0.062 0.050 0.050 0.054
4 0 0.050 0.055 0.047 0.052 0.052
0 0.4 0.191 0.217 0.189 0.216 0.065
4 0.4 0.064 0.220 0.063 0.217 0.048
The average distance between two nodes in real-life networks is low (Amaral et al. (2000)). To obtain a graph with a high degree of clustering but a low average path length, one can rewire the edges of a $K$-regular graph with some probability, say $p$. This parametrization characterizes a small world network. The higher the rewiring probability, the lower the average path length. Another class of random graph models that achieves the same property of a high clustering coefficient but a low average path length is that of the stochastic block model. Consider a graph of size $N$ with a partition of nodes. The size of the parts are $\{n_{1},n_{2},...,n_{K}\}$. These parts represent clusters in the graph, and we can model a random graph using it by defining the probability of an edge between nodes within a part and between two parts. Hence, we obtain a probability (symmetric) matrix of size $K\times K$. Note that because we want a high degree of clustering in the graph, the diagonal elements of the probability matrix are typically higher than the off-diagonal elements. The probability matrix is also called the preference matrix in the literature. We now define the setup for the simulation study:
@@
    Direct treatment effect: We take $\tau_{direct}$ equal to 0 and 4 in our simulations.
@@
    Test statistic: We consider the $T_{quant}$ test statistic described in Section 4.3. It takes the difference in the treatment effect of control units in the first quartile of the proportion of treated neighbors to control units in the last quartile. We define the same for treated units and take the weighted average. We also consider the $T_{bond}$ test statistic, proposed in Bond et al. (2012), defined as follows:
    $T_{bond}(G|Z_{obs})=\frac{\sum_{i,j\in[N]}G_{ij}\cdot Z_{j}\cdot Y_{i}}{\sum_{i,j\in[N]}G_{ij}\cdot Z_{j}}-\frac{\sum_{i,j\in[N]}G_{ij}\cdot(1-Z_{j})\cdot Y_{i}}{\sum_{i,j\in[N]}G_{ij}\cdot(1-Z_{j})}.$
The null distribution was generated by sampling 1000 graphs with the same degree sequence as the observed graph and isomorphic to the observed graph in a second option. The p-value is calculated as the sample proportion of times the imputed test statistic for the sampled graph realizations is more than the observed test statistic. We reject the null hypothesis if the p-value obtained is less than the chosen significance level of 0.05. We replicated the above-described setup 4000 times and calculated the proportion of times we rejected the null. We present this result, namely rejection rate, in the Table 1. For robustness, we perform the simulations for two network-generating processes. We obtain the rejection rates for the same setup for the procedure proposed by Athey et al. (2018). We implement random focal vertex selection and show our results for the Edge-Level-Contrast test statistic (Athey et al. (2018)). We label this test as ‘Conditional Focal Test’.
We then perform the simulations for cluster randomized experimental setup. We present the results in Table 2. The total number of units in the simulation is the same as before, that is, 599. We obtain the clusters in the graph by implementing the epsilon-net clustering algorithm, described in Algorithm 1 (Appendix A). We take epsilon to be equal to 3. This implies all the clusters contain units less than three hops away from each other, whereas the between-cluster distance is at least three hops for any two units. We take the probability of treating a cluster to be 0.5. We implemented a Bernoulli trial at the cluster level. Under Procedure 3, the null distribution is approximated by taking 1000 samples of graphs isomorphic to the original graph, such that the corresponding subgraphs on treated units and controlled units remain isomorphic. This is obtained by considering permutations where treated units can be mapped only to treated units (and thus, control units can only be mapped to control units). We replicated this setup 4000 times and calculated the proportion of times the proposed test rejected the null. We label Procedure 3 as ‘Block Isomorphism Permutation Test’ in Table 2. This was performed for two network-generating processes as before.
@@
Network Distribution Direct Effect Spillover Effect Block Isomorphism Permutation Test
$T_{bond}$ $T_{quant}$
Small world network 0 0 0.052 0.053
4 0 0.056 0.051
0 0.4 0.429 0.352
4 0.4 0.429 0.348
Stochastic block model 0 0 0.051 0.057
4 0 0.048 0.046
0 0.4 0.225 0.237
4 0.4 0.227 0.242
@@ 5.2 Comparison with related work
We see that our proposed methods have type 1 error controlled appropriately. This corresponds to rows with $\tau_{spill}$ equal to zero in Tables 1 and 2. When $\tau_{spill}$ equals 0.4, our method performs with competitive power compared to existing methods in the literature (Athey et al. (2018)). We highlight the $\tau_{direct}$ equal to 4 and $\tau_{spill}$ equal to 0.4, where our method significantly improves. The significance of the setup lies in its resemblance to many naturally occurring settings, where the direct effect is meaningfully more than the spillover effect (for example, see Cai et al. (2015)). We see in Table 1 that our proposed method boosts power when the clustering of the network-generating process increases, as with the stochastic block model. This contrasts with the method proposed by Athey et al. (2018), where the method does not retain power. In Appendix D, we add more insights on the performance of the $T_{bond}$ test statistic
as observed in Table 1.
An interesting result is in the case of cluster-randomized experiments where a higher degree of clustering shows a drop in power. This can be observed when comparing the power values of the small-world network to the stochastic block model. This can be attributed to a reduction of sample space of imputed values of the test statistic. Hence, a higher degree of clustering increases the cluster sizes, restricting the direct treatment assignment. For instance, an observed graph in the case of the small-world network had 16 clusters, compared to 11 clusters in the case of the stochastic block model.
@@
We also perform the simulation study for cluster randomized experiments for the methods proposed in Athey et al. (2018) and Puelz et al. (2022). As discussed in Section 3, we encounter the degenerate null distribution and do not present the corresponding results. This is more prominent in the method proposed by Athey et al. (2018), compared with Puelz et al. (2022). Since Puelz et al. (2022) considers the experiment’s design when selecting focal units, we see more robustness of the biclique test (Puelz et al. (2022)) where degenerate null distributions are reduced in frequency. Overall, our method stands more robust to degeneracy and performs well in terms of power in our simulations. We also see that the $T_{bond}$ test statistic performs slightly better than $T_{quant}$ test statistic in some cluster-randomized setups. This can be attributed to $T_{bond}$ being weighted by treated numbers in the graph instead of the corresponding quantile-based weights in $T_{quant}$ and the heterogeneity of the clustered assignment setup. We emphasize that the $T_{bond}$ in conjunction with our methodology continues to be valid regardless of the size of the direct effect. The method proposed by Bond et al. (2012) does not work for settings with non-zero direct effects (Refer Athey et al. (2018) for an elaborate discussion).
6 Weather insurance adoption in rural China
@@
We re-analyze a field experiment conducted in rural China on rice crop farmers studying neighborhood interference effects (Cai et al. (2015)). The research motive behind the experiment was to see if social networks among the farmers have any influence on weather insurance adoption. The experiment involved intensive training sessions to promote crop insurance products to farmers against extreme weather. A primary goal of the experiment was to identify if there is information dissemination in the farmers’ community and, if so, what mechanisms drive these information exchanges. These are important questions to be answered in this setting as they are consequential to government policy action. A deeper understanding of insurance product adoption via social network mechanisms helps devise a more cost-effective government policy campaign to enhance product adoption among the target community. An important reason to study interference in such a setup is to avoid biased estimates and accurately evaluate policy performance measures.
@@
We discuss the experimental design setup by the researchers. Researchers hypothesize that access to better information about weather insurance products would increase the adoption of the product in the rice farmers’ community. To capture access to information differentially as treatment and control, two types of information sessions were created: simple sessions of 20 minutes, which provided an introduction to the insurance contract, and intensive sessions of 45 minutes that covered the insurance contract in much more detail, including explaining to them what the benefits are of taking up insurance. To capture information dissemination across the social network, these sessions were conducted in two rounds with a gap of 3 days. The idea behind this setup is to allow the farmers some time to discuss the insurance product among themselves. The interference effect is identified by examining whether households in round 2 with more friends who received intensive sessions in round 1 displayed an enhanced insurance take-up rate. Second-round participants were further randomized into three groups. After the round 2 sessions, the first group was given no additional information, the second group was told about the overall take-up rate in the previous round, and the third group was told in detail about who purchased the product and who did not. This additional information randomization of the second group was done to derive the primary information diffusion mechanism. For illustration purposes, we do not analyze this part of the experiment.
@@
There were 4902 households across 47 different villages in the experiment. A village’s social network is known and assumed to be independent between two distinct villages. Each household is randomized among the 4 different groups: (i) Receiving a simple session in Round 1, (ii) Receiving an intensive session in Round 1, (iii) Receiving a simple session in Round 2, (iv) Receiving an intensive session in Round 2. After the session ends, the households decide to purchase insurance. This is captured as a binary outcome variable of the model, with 1 indicating a purchase decision made and 0 when no purchase is made. Under the assumption that no information diffusion can occur between participant households of Round 1 because of the immediate purchase decision the household had to make, we can assume there are no edges among the household participants in Round 1.
@@
In our analysis, we take all the households receiving simple sessions under the control group and intensive sessions under the treated group. Since spillover can only occur from the treated group of Round 1 to the control group of Round 2, we only consider the network structure amongst these two groups in our study. This includes network structure among households in Round 1 and Round 2 only. Our analysis was run on 433 households of randomly chosen villages in the experiment, of which 222 were in the treated group. We ran our proposed Procedure 2, the degree permutation test, and approximated the null distribution by taking 10,000 samples of graphs with the same degree sequence as the observed graph. We obtained a p-value of 0.0002, which indicates the presence of interference.